OpenStudy (anonymous):

1. y = x + 6 y = 2x (1 point) (6, 12) (–12, –6) (–6, –12) (2, 4) 2. 3x + 2y = 7 –3x + 11 = y (1 point) (6, –3) (6, –7) (-4,19/2) (5, –4) 3. 2x – y = –7 4x – y = –4 (1 point) (–1.5, 4) (1.5, 10) (4, –1.5) (–1.5, –2) 4. –4y + 4 = x 2x + 8y = 8 (1 point) one solution two solutions infinitely many solutions no solution 5. 6x + 2 = y 3y – 18x = 12 (1 point) one solution two solutions infinitely many solutions no solution

OpenStudy (anonymous):

1 (6,12)

OpenStudy (anonymous):

2) (5, -4) 3x + 2y = 7 y = 7/2 - 3x/2 y = 7/2 - (3(5)/2) y = 7/2 -15/2 y = -8/2 y = -4 and (6, -7) –3x + 11 = y y = 11-3x y = 11 - 3(6) y = 11 - 18 y = -7